1. Use the Z table to find the probability associated with each of the following areas...

1.

Use the Z table to find the probability associated with each of the following areas under a normal curve:

a. Above a z-score of 1.67

b.Above a z-score of–1.35
Answer #1

Solution:

a)

Area above a z-score of 1.67

= P(Z > 1.67)

= 1 - P(Z < 1.67)

= 1 - 0.9525

= 0.0475

b)

Area above a z-score of–1.35

= P(Z > -1.35)

= 1 - P(Z < -1.35)

= 1 - 0.0885

= 0.9115